### v1

# Attention-Based Beamformer For Multi-Channel Speech Enhancement

Abstract
Minimum Variance Distortionless Response (MVDR) is a classical adaptive beamformer that theoretically ensures the distortionless transmission of signals in the target direction. Its performance in noise reduction actually depends on the accuracy of the noise spatial covariance matrix (SCM) estimate. Although recent deep learning has shown remarkable performance in multi-channel speech enhancement, the property of distortionless response still makes MVDR highly popular in real applications. In this paper, we propose an attention-based mechanism to calculate the speech and noise SCM and then apply MVDR to obtain the enhanced speech. Moreover, a deep learning architecture using the inplace convolution operator and frequency-independent LSTM has proven effective in facilitating SCM estimation. The model is optimized in an end-to-end manner. Experimental results indicate that the proposed method is extremely effective in tracking moving or stationary speakers under non-causal and causal conditions, outperforming other baselines. It is worth mentioning that our model has only 0.35 million parameters, making it easy to be deployed on edge devices.
Index Terms:
MVDR, attention-based mechanism, multi-channel speech enhancement, inplace convolution
00footnotetext: * These authors contributed equally to this work.
I Introduction
For microphone array signal processing [1, 2], most speech enhancement methods are based on fixed and adaptive beamformer techniques [3, 4], leveraging the spatial characteristics of multi-channel data. Minimum variance distortionless response (MVDR) [4] uses spatial information to calculate beamforming coefficients, ensuring that signals from the desired direction are transmitted without distortion.
Recently, mask-based beamforming approaches [5, 6, 7, 8] have attracted increased attention. The mask-based beamformer utilizes time-frequency masks derived from neural networks [5, 6] or other models [7, 8] to compute spatial covariance matrices (SCMs), which capture the spatial information of the sources. These techniques are adaptable to both time-invariant and time-varying environments. For time-invariant systems, the methodology involves computing the SCM through temporal averaging, ensuring stability over periods without significant change. Time-varying systems necessitate a dynamic approach to SCM estimation, employing either recursive methods[7] or block processing [8] to adapt to real-time changes. This dynamic adaptation faces the critical challenge of selecting an appropriate forgetting factor for recursive updates or determining the optimal block size for block processing.
ATT-MVDR [9] proposes a two-stage approach to address the above problems. In the first stage, the mask is estimated using a neural network with the fifth channel as the target. In the second stage, two Transformer [10] encoder architectures are employed to allocate the temporal weights for the speech and noise instantaneous SCMs (ISCMs) [9] estimated by the mask. However, using the mask from the fifth channel for the SCMs calculated by all channels can easily result in error accumulation. Moreover, the Transformer encoder is computationally intensive and parameter-heavy, limiting its deployment in edge devices.
The traditional beamforming shows that the spatial information exists in each frequency bin of the array signal spectrum [11]. Wideband beamforming is also independently processed in each frequency bin, which we call an inplace process. When ATT-MVDR uses Transformer encoders to estimate attention weights, the frequency dimension is reduced through the linear layer, complicating spatial information learning. Recently, the Inplace Gated Convolutional Recurrent Neural Network (IGCRN) [12] is proposed to learn spatial information across channels by discarding the down-sampling operation in the frequency dimension from the Convolutional Recurrent Network (CRN) [13][14]. Due to the inplace characteristic of IGCRN, the spatial cues are explicitly maintained within each frequency bin. In addition, the LSTM layer is designed to capture temporal correlations across channels for each frequency bin. Since the time delay for a specific direction is similar among different frequency bins, they share the same LSTM blocks, referred to as the frequency-independent LSTM. The IGCRN has been successfully applied to mono acoustic echo cancellation (AEC)[15], and stereo AEC[16].
Figure 1: (a) An overview of ABIC-MVDR. (b) Structure of encoder and decoder. (c) Inplace self-attention module (ISAM).
This paper introduces a novel attention-based inplace convolution recurrent neural network combined with a beamformer for multi-channel speech enhancement, referred to as ABIC-MVDR. Firstly, masks for the speech and noise ISCMs are predicted using the IGCRN. The inplace convolution operator and two-layer frequency-independent LSTM are employed to extract spatial information and analyze inter-channel time delays efficiently. To better adapt to different environments, we propose an innovative approach that applies two decoder outputs to estimate the temporal attention weights for SCM reconstruction, similar to the self-attention mechanism. By optimizing the model in an end-to-end manner, the accumulation of errors due to multiple stages can be avoided. Our method outperforms baselines with reduced computation and fewer parameters.
II Methodology
II-A Problem definition
Let $S^{m}_{t,f}$ and $N^{m}_{t,f}$, where $t$ and $f$ represent the time frame index and frequency bin index respectively, denote the clean speech and background noise received by microphone $m$. The received signal, denoted as $Y^{m}_{t,f}$, can be modeled as:
$Y^{m}_{t,f}$ $=S^{m}_{t,f}+N^{m}_{t,f}$ (1)
$=S_{t,f}\cdot H^{m}_{f}+N^{m}_{t,f}$
Here, $H^{m}_{f}$ indicates the acoustic impulse response from the target speaker to the $m$-th microphone.
II-B An overview of ABIC-MVDR
The proposed ABIC-MVDR system comprises three primary components: the IGCRN backbone network (IBN), the Inplace self-attention module (ISAM), and the MVDR beamformer. As shown in Fig. 1 (a), the signal collected by multiple microphones is processed through Short-Time Fourier Transform (STFT). The real and imaginary parts are concatenated with the channel dimension and fed into the encoder. The IBN is adopted to predict masks for the speech and noise ISCMs. Two ISAMs are then utilized to reassign weights over time for the speech and noise SCMs reconstruction. Finally, the MVDR beamformer is applied to reconstruct the target speech.
II-C IGCRN backbone network (IBN)
As depicted in Fig. 1 (a), the model comprises one encoder and five decoders. Five decoders share the similar structure: 1) Two decoders are used to build the self-attention mechanism in speech ISAM; 2) Two decoders are used to build the self-attention mechanism in noise ISAM; 3) The remaining decoder prediction masking is used to build the ISCM.
The encoder and decoder are shown in Fig. 1 (b). Inplace-convolution refers to a convolutional neural network where the stride of the kernel is set to one, ensuring that features are not downsampled in the frequency dimension. This approach naturally and explicitly preserves spatial correlations within each frequency bin. Skip connections are used to concatenate the output of each Inplace-Conv2d block to the input of the corresponding Inplace-Deconv2d block. Following each Inplace-Conv2d and Inplace-Deconv2d block, a batch normalization[17] and an exponential linear unit (ELU) activation function are sequentially applied. A two-layer LSTM is utilized to model the temporal characteristics of each frequency bin across channel, effectively analyzing spatial information through time delay. Since the time delay for a specific direction is consistent across different frequency bins, they share the same LSTM units. Following this, a channel-wise linear layer is implemented to ensure that the output shape of the LSTM matches the input.
II-D Inplace Self-Attention Module (ISAM)
Unlike the transformer encoder architecture used in [9], we compute attention weights for the ISCM through a transposed matrix multiplication between the outputs from two IGCRN decoders in ISAM. In Fig. 1 (a), two ISAMs are utilized to allocate weights over time for reconstructing the speech SCM, denoted as $\Phi^{S}_{t,f}$, and the noise SCM, denoted as $\Phi^{N}_{t,f}$, respectively. The details of the ISAM are shown in Fig. 1 (c). The decoder structure in ISAM is similar to that described in II-C. The $\tanh$ activation function is employed in the final layer. The two decoders in ISAM predict the $\mathbf{query}$ and $\mathbf{key}$ vectors, respectively.
After obtaining $\mathbf{query}\in\mathbb{R}^{F\times T\times D}$ and $\mathbf{key}\in\mathbb{R}^{F\times T\times D}$, the attention weights $\mathbf{w}^{\text{att}}\in\mathbb{R}^{F\times T\times T}$ are calculated by (2), where the softmax is applied on the second $T$ dimension. Here, $T$ and $F$ represent the time frame index and frequency bin index, respectively, and $D$ represents the feature dimension, set to 24 in this paper.
$\mathbf{w}^{\text{att}}=\text{softmax}\left(\text{mask}\left(\frac{\mathbf{query}(\mathbf{key})^{T}}{\sqrt{D}}\right)\right)$ (2)
The mask is a lower triangular matrix to ensure the self-attention mechanism is causal, as in (3). If the model is operating under non-causal conditions, the mask function should be omitted.
$\text{mask}(x(f,i,j))=\begin{cases}x(f,i,j),&\text{if }i\geq j\\ -\infty,&\text{otherwise}\end{cases}$ (3)
The SCM $\mathbf{\hat{\Phi}}\in\mathbb{C}^{F\times T\times M\times M}$ is calculated using $\mathbf{w}^{\text{att}}$ and the $\mathbf{value}$ where M represents channels, following (4).
$\mathbf{\hat{\Phi}}=\mathbf{w}^{\text{att}}\cdot\mathbf{value}$ (4)
Here, the $\mathbf{value}$ refers to the estimated ISCM obtained after the ISCM Masking Process. $\mathbf{\hat{\Phi}}$ can represent either the speech SCM or the noise SCM, depending on the corresponding ISCM and $\mathbf{w}^{\text{att}}$.
II-E Minimum Variance Distortionless Response Beamformer
In this paper, the beamforming filter coefficients $\mathbf{w}_{t,f}\in\mathbb{C}^{M}$ [18] are calculated using MVDR, as shown in (5). The $\boldsymbol{\Phi}^{S}_{t,f}\in\mathbb{C}^{M\times M}$ and $\boldsymbol{\Phi}^{N}_{t,f}\in\mathbb{C}^{M\times M}$ are the estimated SCMs of the speech and noise signals with $M$ channels at time-frequency bin $(t,f)$, respectively. The $\mathbf{u}\in\mathbb{C}^{M}$ represents the index of the reference microphone.
$\begin{split}\mathbf{w}_{t,f}=\frac{(\boldsymbol{\Phi}^{N}_{t,f})^{-1}\boldsymbol{\Phi}^{S}_{t,f}}{\mathrm{Tr}((\boldsymbol{\Phi}^{N}_{t,f})^{-1}\boldsymbol{\Phi}^{S}_{t,f})}\mathbf{u},\end{split}$ (5)
The enhanced speech is calculated through the observed signal $\mathbf{Y}_{t,f}$, as shown in (6). Here, $(\cdot)^{H}$ represents the conjugate transpose.
$\begin{split}\hat{\mathbf{S}}_{t,f}=\mathbf{w}^{H}_{t,f}\mathbf{Y}_{t,f},\end{split}$ (6)
III Experiments
III-A Dataset and setting
We generate the scenario of moving speakers in noisy conditions for training and testing. The speech signals are sourced from the WSJ0 corpus[19], while the noise signals are obtained from the CHiME-3 corpus[20]. For ease of comparison with [9], we design the experimental scenario to match that of [9], with the only difference being that speech and noise signals are randomly paired at diverse signal-to-noise ratios (SNRs) ranging from -10 dB to 10 dB. After statistical analysis of the generated moving source signals, the minimum, maximum, and average source velocities $[m/s]$ (moving distance $[m]$/utterance duration $[s]$) are $0.012\ m/s$, $2.404\ m/s$, and $0.238\ m/s$, respectively. We also create a non-moving dataset to evaluate scenarios with fixed-position speakers. Its configuration is identical to that of the moving dataset, with the sole difference being that the source speaker’s position is fixed at the starting location. Through the above approach, we produce 30,000 noisy speech signals for the training set and 2,000 for the development and evaluation sets in both the non-moving and moving datasets, respectively.
TABLE I: The proposed ABIC-MVDR architecture, where B, T and F represent batch size, frame, and frequency bin, respectively.
Layer name Input size Hyperparameters Output size
Inplace Conv2d (1) [B, 10, F, T] $5\times 1,(1,1),24$ [B, 24, F, T]
Inplace Conv2d (2-6) [B, 24, F, T] $5\times 1,(1,1),24$ [B, 24, F, T]
Reshape [B, 24, F, T] - $[B\times F,T,24]$
Channel-wise LSTM×$\times 2$ $[B\times F,T,24]$ 48 $[B\times F,T,48]$
Channel-wise Linear $[B\times F,T,48]$ 24 $[B\times F,T,24]$
Reshape $[B\times F,T,24]$ [B, 24, F, T]
Inplace Deconv2d (6-2) ×$\times 5$ [B, 48, F, T] $5\times 1,(1,1),24$ [B, 24, F, T]
Inplace Deconv2d (1) (IBN) [B, 48, F, T] $5\times 1,(1,1),24$ [B, 1, F, T]
Inplace Deconv2d (1) ×$\times 4$ (ISAM) [B, 48, F, T] $5\times 1,(1,1),24$ [B, 24, F, T]
III-B Model configuration
[…]
TABLE II: STOI[%], ESTOI[%], PESQ, SI-SDR, WER[%] and TSOS[%] for Non-Moving and Moving Datasets.
Model Non-Moving Dataset Moving Dataset
STOI $\uparrow$ ESTOI $\uparrow$ PESQ $\uparrow$ SI-SDR $\uparrow$ WER $\downarrow$ TSOS $\downarrow$ STOI $\uparrow$ ESTOI $\uparrow$ PESQ $\uparrow$ SI-SDR $\uparrow$ WER $\downarrow$ TSOS $\downarrow$
Noisy 75.39 55.03 2.12 2.18 41.87 - 74.93 54.71 2.12 2.09 45.94 -
Non-causal F-Conv-TasNet 84.18 69.74 2.46 10.12 49.46 21.74 83.99 69.35 2.48 10.05 50.63 19.63
ATT-MVDR 87.75 73.69 2.69 10.49 30.43 0.86 85.47 70.57 2.63 9.78 38.01 0.44
ABIC-MVDR 92.41 81.62 3.05 12.69 18.63 0.52 91.49 79.98 3.03 12.47 22.02 0.76
Causal F-Conv-TasNet 83.09 68.03 2.41 9.69 50.20 17.12 82.66 67.49 2.42 9.66 52.98 17.58
ATT-MVDR 86.3 71.19 2.58 9.03 32.37 0.49 83.8 67.75 2.52 8.57 40.38 0.51
ABIC-MVDR 91.88 80.62 2.99 11.22 19.82 0.42 90.46 77.99 2.92 11.54 28.17 0.43
IV Experiment results and analysis
IV-A Experimental results
In this paper, the ATT-MVDR [9] is employed as the baseline. The output of the first stage in ATT-MVDR is referred to as F-Conv-TasNet [23], which a single-channel model used to estimate the target speech mask, as described in I. Table II compares our method with baselines across six metrics under different combinations of conditions (moving or non-moving and causal or non-causal). Table IV shows the parameters and Multiply-Accumulate Operations (MACs) for each model, where the MACs are calculated for 1 second audio. Six objective metrics are employed: short-time objective intelligibility (STOI) [24], extended short-time objective intelligibility (ESTOI) [25], perceptual evaluation of speech quality (PESQ) [26], scale-invariant signal-to-distortion ratio (SI-SDR) [22], word error rate (WER), and target speaker over-suppression (TSOS)[27]. To evaluate the degree of speech distortion, WER and TSOS are employed as key metrics, with lower values indicating less distortion. WER is computed using the Whisper model with its base size [^1] https://github.com/openai/whisper.
F-conv-tasnet estimates the T-F mask using a neural network, and its WER and TSOS are significantly higher than that of models using MVDR, such as ABIC-MVDR and ATT-MVDR, which shows that one of the advantages of MVDR is that it reduces speech distortion. As shown in Table II and Table IV, the second stage in ATT-MVDR has limited improvement over the first stage, but consumes a lot of parameters and computing resources. As discussed in I, one explanation is the error accumulation across the two stages and the lack of consideration for frequency band independence when calculating attention weights. Our proposed method outperforms other baselines across most metrics under different conditions, with the exception of TSOS, which is sometimes slightly worse. ABIC-MVDR has only 0.35 million trainable parameters, which is 108 times fewer than ATT-MVDR, and requires 4.04 G/s MACs, approximately 1.68 times fewer. This demonstrates our IGCRN-based method is more effective in learning spatial information and calculating attention weights for SCM reconstruction compared to ATT-MVDR.
IV-B Ablation experiments
We conduct 3 ablation experiments to verify the effectiveness of each module:
  1. 1.
    ONLINE-MVDR / BLOCK-MVDR: To validate the facilitation of the ISAM module, this ablation experiment calculates the SCM using Adaptive-based [7] (ONLINE-MVDR) and Block-based [8] (BLOCK-MVDR) based on the ISCM. The forgetting factor is set to 0.995 in ONLINE-MVDR and the block size is set to 30 in BLOCK-MVDR.
  2. 2.
    CRN-MVDR: The IGCRN module in the proposed method is replaced by the CRN[28] to verify the performance of the IGCRN.
TABLE III: Ablation experiments of ABIC-MVDR on moving datasets.
model STOI $\uparrow$ ESTOI $\uparrow$ PESQ $\uparrow$ SI-SDR $\uparrow$ WER $\downarrow$ TSOS $\downarrow$
Noisy 74.93 54.71 2.12 2.09 45.94 -
non-causal CRN-MVDR 90.15 77.82 2.91 11.89 26.60 0.62
ABIC-MVDR 91.49 79.98 3.03 12.47 22.02 0.76
causal ONLINE-MVDR 82.17 64.12 2.37 2.12 43.75 4.63
BLOCK-MVDR 82.36 63.39 2.33 3.97 42.77 1.72
CRN-MVDR 89.90 77.20 2.84 11.28 26.89 1.00
ABIC-MVDR 90.46 77.99 2.92 11.54 28.17 0.43
Table III shows the results of the ablation experiments on moving datasets. Compared to ONLINE-MVDR and BLOCK-MVDR, the proposed model demonstrates superior performance. This suggests that the inclusion of the ISAM module enhances the capability of the SCM reconstruction and target speaker tracking in complex scenarios. From Table III and IV, it can be seen that the proposed model outperforms CRN-MVDR, with reduced computation and fewer parameters. This proves that spatial information is effectively extracted through IGCRN for reconstructing SCMs, making it compatible with beamformer.
TABLE IV: Comparison of Model Parameters and MACs
Method Parameters (M) MACs (G/s)
F-conv-Tasnet 8.81 2.43
ATT-MVDR 37.99 6.77
CRN-MVDR 19.82 6.04
ABIC-MVDR 0.35 4.04
V Conclusion
This paper introduces a novel approach that combines IGCRN and an attention-based mechanism to estimate speech and noise SCM, followed by applying MVDR to obtain enhanced speech. The IGCRN leverages the characteristics of the inplace convolution operator and frequency-independent LSTM to learn spatial information. The attention-based mechanism is applied to allocate temporal weights for reconstructing the speech and noise SCMs accurately. The experimental results demonstrate that the proposed method performs baselines under various conditions. Furthermore, due to the small number of parameters, the proposed model can be easily deployed in edge devices.